# An Efficient TLS 1.3 Handshake Protocol with VC Certificate Type

### v1

@@ Abstract
The paper presents a step forward in the design and implementation of a Transport Layer Security (TLS) handshake protocol that enables the use of Verifiable Credential (VC) while maintaining full compliance with RFC-8446 and preserving all the security features of TLS 1.3. The improvement over our previous work lies in the handshake design, which now only uses messages already defined for TLS 1.3. The design has an incredibly positive impact on the implementation, as we made minimal changes to the OpenSSL library and relied mostly on a novel external provider to handle VC and Decentralized IDentifier related operations. The experimental results prove the feasibility of the design and show comparable performance to the original solution based on PKI and X.509 certificates. These results pave the way for the adoption of Self-Sovereign Identity in large-scale IoT systems, with a clear benefit in terms of reducing the cost of identity management.
@@ I Introduction
The number of Internet of Things (IoT) devices is multiplying every year. According to Statista [1], the number of IoT nodes is expected to grow to 29 billion by 2030. At the same time, IoT systems, with their poor security reputation, have become an entry point for adversaries to target critical IoT systems. All leading cybersecurity analysts report a steady increase in attacks against IoT nodes from year to year. Providing these devices with secure digital identities is a critical first step in keeping pace with evolving threats. Especially in the case of large-scale IoT systems, automation is crucial for provisioning, updating and revoking node identities. Current solutions rely on a centralized PKI with X.509 certificates [2]. However, this process is largely seen by industry partners as inefficient and costly. The use of PKI and X.509 certificates requires human intervention as there is no standard automated solution tailored to IoT nodes and protocols. For instance, anytime the identity of a node changes or its public key is rotated, the X.509 certificate must be revoked and a new one must be issued. The revoked certificate must be added to a certificate revocation list, which is not immediately propagated throughout the IoT system. On top of this, the Certificate Authorities comes with other drawbacks. They impose a cost on the system administrator to add a trusted third party into trust relationships, and they are vulnerable to a single point of failure, as a single breach can compromise all issued digital certificates, requiring an update of all node identities before network operations can resume.  [3, 4, 5, 6, 7].
@@
The Self-Sovereign Identity (SSI) model [8] represents an alternative solution to overcome the limitations of the centralized PKI and reduce the complexity and the related cost of X.509 certificate management in large scale IoT systems. SSI is an emerging decentralised digital identity model. It gives a node full control over the data it uses to generate and prove its identity. An SSI native IoT node generates its identity key pair and stores the public key in a distributed ledger for other nodes to authenticate it. A node’s Decentralized IDentifier (DID) represents the address in the distributed ledger where other nodes can retrieve its public key. Once these two components have been generated, a node can request a Verifiable Credential (VC) from the Issuer in the IoT system. The VC contains the metadata and claims about the identity of the node that holds it. The combination of the key pair, the DID and at least one VC shapes the digital identity in an SSI native IoT system. The composition of the digital identity reflects the decentralised nature of the SSI model. In fact, the distributed ledger overcomes the single point of failure limitation of the traditional PKI as no single party is in control of the ledger that contains all public keys. Plus, the separation of public key information (stored in the ledger) from identity attributes (stored in VCs) also comes with several advantages: (i) a node can rotate its key pair without the need of refreshing the VC, by updating autonomously the information in the ledger, (ii) a node can immediately revoke its DID if the key pair gets compromised invalidating right away the VC, (iii) the identity life cycle of a device does not require human intervention, and (iv) the scale at which identity can be provided to devices increases dramatically.
@@
In light of these advantages our objective is to contribute to the adoption of SSI in IoT systems and its use for authentication purposes. However, most of the discussions and proposals for the adoption of the SSI model consider an implementation at the application layer of the TCP/IP stack [9, 10]. The problem with application layer authentication is, first, the lack of established protocols that can be used in IoT systems and, second, the need to use two different technologies simultaneously, X.509 certificates on the server side to establish a secure channel and VCs on the client side to complete the mutual authentication on top of the established secure channel. This creates an overall friction in the adoption of the SSI model.
@@
We believe in implementing (mutual) authentication at the transport layer of the TCP/IP stack directly within the Transport Layer Security (TLS) protocol [11] with VC. Using VC in TLS maximises the benefits of adopting SSI while avoiding the use of PKI and X.509 certificates and their associated costs. In addition, performing authentication at a lower layer in the stack reduces the attack surface and avoids potential vulnerabilities that the application layer may add.
@@
Enabling the use of VCs requires extending the TLS 1.3 handshake protocol to work with VCs in addition to the existing certificate types (e.g., X.509, raw public key) to ensure interoperability. We have already proposed an extension of the TLS 1.3 handshake protocol in [12] enabling the use of VCs and DIDs while preserving all security features of the original handshake. The design has been validated with a real implementation in OpenSSL and statistically relevant experimental results in real-world settings. The design in [12] enables the use of VC and DID for authentication by adding an extension and two new messages fully compliant with RFC-8446 [11]. Despite full compliance and interoperability our design required a modification to the state machine of the protocol implementation. There is an high cost associated to the modification and associated security review of the TLS handshake state machine. This paper present a new efficient design of the TLS 1.3 handshake protocol that minimizes the changes with respect to the original one hence make it more appealing for a possible standardization [13] and upstream contribution to the OpenSSL library.
@@
While working on the design and implementation of the new efficient handshake protocol, an interesting work appeared in literature [14]. The authors acknowledge our previous work [12, 15] and they have the same goal of performing authentication in the TLS with VCs and DIDs. Their design divides the overall authentication process in two phases: the pseudo-anonymous authentication with DIDs in the TLS handshake protocol and the identification with Verifiable Presentations (VPs) in the TLS record protocol. However, to leave the handshake unchanged they wrap DIDs inside self-signed X.509 certificates, but this implies the overhead of the certificate generation and management, and the exchange of unnecessary data. In addition, their overall (mutual) authentication process requires a higher number of RTTs than our solutions, because of the identification phase. Furthermore, the pseudo-anonymous authentication with DIDs establishes the secure channel even though the identity of the endpoints is still unverified. In a mutual authentication scenario any adversary can create its own DID and establish the secure channel with the only purpose of draining resources on server side. In our solutions an endpoint never finalizes the handshakes with an unidentified endpoint.
@@
This paper present our new design of an efficient TLS handshake protocol. It does not add any new message or delete any existing one from the original handshake, but only embeds a new extension and adds the VC certificate type to the list of accepted values in the client_certificate_type and server_certificate_type defined in [16]. These two extensions provide by default hybrid handshakes and a fallback mechanism to X.509 certificates in case other certificate types are not supported. This novel approach leaves untouched the core of the handshake protocol and it is meant to ease adoption. We implemented most of the additional functionalities in an external provider for OpenSSL.
@@ II Self-Sovereign Identity
Layer 1 is implemented by any Distributed Ledger Technology (DLT) acting as a RoT for public identity data. DLTs are distributed and immutable means of storage by design [17]. The Decentralized IDentifier (DID) [18] is the new type of globally unique identifier designed to verify a node. The DID is a Uniform Resource Identifier (URI) in the form
@@
While Layer 2 uses DIDs to start authentication, Layer 3 completes it and also deals with authorisation to services/resources using Verifiable Credentials (VCs) [20]. A VC is a digital credential that contains additional characteristics of a node’s identity beyond its identity key pair, the DID, and the DID Document.
@@
  • Holder is the node that owns one or more VCs and generates a Verifiable Presentation (VP) to request access to services/resources from a Verifier. An Holder builds the VP as an envelope of the VC and signs it with its identity private key $sk$;
@@ III-A Original handshake
The TLS 1.3 handshake protocol establishes the secure channel by exchanging the messages shown in Fig. 1(a). Upon the handshake, client and server negotiate cryptographic parameters through the exchange of ClientHello and ServerHello messages. Those parameters are the symmetric cipher and hash algorithm used to ensure confidentiality and integrity respectively. The client and the server generate a secret by using ephemeral Diffie–Hellman(DH) key exchange and use the HMAC-based Extract-and-Expand Key Derivation Function (HKDF) algorithm with the negotiated hash algorithm to derive the session keys from the secret. Then, the server authenticates with the client by sending the Certificate message containing the certificate chain (Root CA certificate excluded), and the CertificateVerify message which is the signature over all the exchanged messages computed with its private key. The client verifies the validity of the certificate chain and the signature in the CertificateVerify message to check the identity of the server. The server can request client authentication by sending the CertificateRequest message. The client authenticates with the server in the same way.
@@ III-B Handshake with VC
We achieve these objectives by leveraging the client_certificate_type and server_certificate_type extensions introduced by RFC-7250 [16] in the TLS 1.3 handshake protocol. These two extensions are used to negotiate certificate types different than X.509 certificates.
We added the VC certificate type to the already existing X.509 and RawPublicKey types to enable the use of VCs. The exchange of the messages in the TLS handshake with the new certificate type is represented in Fig.1(b). The client sends the two extensions in the ClientHello indicating the support for the VC certificate type for both client and server authentication. The server sends the two extensions in the EncryptedExtensions message selecting the VC type from the list presented by the client for both the client and the server.
@@
The did_methods is sent in the ClientHello and EncryptedExtensions messages. A client that sends the server_certificate_type extension containing the VC certificate type must also send the did_methods extension. Similarly a server that selects VC as the certificate type for the client_certificate_type extension must also send the did_methods extension. A server must send a list of DID Methods both client and server share if it receives the did_methods extension from the client.
@@
The use of VCs reduces the total number of signatures in the handshake. VC Issuers are not organized hierarchically. The server must verify only the signature of the Issuer in the Certificate message. When using X.509 certificates it must validate the certificate chain up to the Root CA. This is considered an advantage in the case of resource constrained devices as it can improve performance while maintaining compliance with the VP concept.
Upon receiving CertificateVerify message, the client checks that the VC follows the scheme specified in the @context field, checks the validity of the VC metadata, resolves the DID of the Issuer and verifies its signature on the VC, and then extracts the server DID from the credentialSubject field of the VC and resolves the server DID to retrieve the server public key from the ledger. Finally, the client verifies the signature in the CertificateVerify message. Note that in the context of IoT systems, the number of available Issuer(s) is expected to be limited (one very often), therefore both client and server could securely maintain the public key of Issuer(s) they trust to avoid the DID resolution. The client authenticates with the server in the same way.
@@ IV-A Security
The TLS handshake protocol with VC certificate type maintains the same security features of the original handshake as the design remains the same. The only addition is the did_methods extension. When the extension is sent in the ClientHello message it benefits from the integrity and authentication properties. When it is sent in the CertificateRequest message it benefits also from confidentiality.
@@
Most of the security discussion on our design focuses on the interaction with the DLT node for peer authentication. In the case of a DID resolution performed in clear, an active man-in-the-middle (MiTM) could impersonate the DLT node, forge a DID document containing the authenticating endpoint’s DID, associate it with a key pair that he owns, and then return it to the DID resolver. Thus, the attacker is able to compute a valid CertificateVerify message by possessing the long term private key. In practice, the MiTM attacker breaks in transit the immutability feature of the DLT (i.e., the RoT for identity public keys). To mitigate this risk an endpoint can establish a secure channel towards the DLT node, either through the Internet Protocol Security (IPSec) [21, 22, 23] or TLS. They both provide authentication of the node and integrity and confidentiality of the data exchanged but they come with an additional overhead. However, there are some differences among the two approaches.
@@
An IPSec channel works at network layer and it is established before the TLS handshake takes place and can be reused until it expires. The moment the endpoint performs the DID Resolve the messages are immediately ciphered and exchanged. An IPSec channel can only be established in case of full control of a DLT node, because additional software must be installed on it. In the case of IoT systems we consider the DLT nodes to be internal to the system and so under full control of the administrator. IPSec employs the Internet Key Exchange (IKE) [22] protocol to authenticate the parties. In IKE authentication must be mutual. One option could be the usage of X.509 certificates, but it is not convenient because all the devices in the system should have an X.509 certificate and so they can directly run the original TLS 1.3 handshake protocol making VCs worthless. Raw Public Keys (RPKs) [16] represents a better option for authentication in this case. Every IoT node must save locally the public key of each DLT node and each DLT node must save locally the key of each IoT node. The number of DLT nodes within an IoT system is expected to be very low (i.e., one or a couple of nodes) with respect to the total number of IoT and edge nodes in the system, so the number of public keys stored by the latter is minimal. Only the DLT nodes should have the full list of public keys of the devices in the IoT system. One could argue that devices in the IoT system could communicate securely by using directly RPKs for authentication in the TLS handshake, but this means that each IoT and edge node should store locally the public keys of all the other nodes in the system and this is again critical in terms of cost associated to identity management.

### v2

@@ Abstract
The paper presents a step forward in the design and implementation of a Transport Layer Security (TLS) handshake protocol that enables the use of Verifiable Credential (VC) while maintaining full compliance with RFC-8446 and preserving all the security features of TLS 1.3. The improvement over our previous work lies in the handshake design, which now only uses messages already defined for TLS 1.3. The design has an incredibly positive impact on the implementation, as we made minimal changes to the OpenSSL library and relied mostly on a novel external provider to handle VC and Decentralized IDentifier (DID) related operations. The experimental results prove the feasibility of the design and show comparable performance to the original solution based on Public Key Infrastructure (PKI) and X.509 certificates. These results pave the way for the adoption of Self-Sovereign Identity (SSI) in large-scale Internet of Things (IoT) systems, with a clear benefit in terms of reducing the cost of identity management.
@@ I Introduction
The number of Internet of Things (IoT) devices is multiplying every year. According to Statista [1], the number of IoT nodes is expected to grow to 29 billion by 2030. At the same time, IoT systems, with their poor security reputation, have become an entry point for adversaries to target critical IoT systems. All leading cybersecurity analysts report a steady increase in attacks against IoT nodes from year to year. Providing these devices with secure digital identities is a critical first step in keeping pace with evolving threats. Especially in the case of large-scale IoT systems, automation is crucial for provisioning, updating and revoking node identities. Current solutions rely on a centralized Public Key Infrastructure (PKI) with X.509 certificates [2]. However, this process is largely seen by industry partners as inefficient and costly. The use of PKI and X.509 certificates requires human intervention, as there is no standard automated solution tailored to IoT nodes and protocols. For instance, anytime the identity of a node changes or its public key is rotated, the X.509 certificate must be revoked, and a new one must be issued. The revoked certificate must be added to a certificate revocation list, which is not immediately propagated throughout the IoT system. On top of this, the Certificate Authorities (CAs) come with other drawbacks. They impose a cost on the system administrator to add a trusted third party into trust relationships, and they are vulnerable to a single point of failure, as a single breach can compromise all issued certificates, requiring an update of all node identities before network operations can resume.  [3, 4, 5, 6, 7].
@@
The Self-Sovereign Identity (SSI) model [8] represents an alternative solution to overcome the limitations of the centralized PKI and reduce the complexity and the related cost of X.509 certificate management in large scale IoT systems. SSI is an emerging decentralized digital identity model. It gives a node full control over the data it uses to generate and prove its identity. An SSI native IoT node generates its identity key pair and stores the public key in a distributed ledger for other nodes to authenticate it. A node’s Decentralized IDentifier (DID) represents the address in the distributed ledger where other nodes can retrieve its public key. Once these two components have been generated, a node can request a Verifiable Credential (VC) from the Issuer in the IoT system. The VC contains the metadata and claims about the identity of the node that holds it. The combination of the key pair, the DID and at least one VC shapes the digital identity in an SSI native IoT system. The composition of the digital identity reflects the decentralized nature of the SSI model. In fact, the distributed ledger overcomes the single point of failure limitation of the traditional PKI, as no single party is in control of the ledger that contains all public keys. Plus, the separation of public key information (stored in the ledger) from identity attributes (stored in VCs) also comes with several advantages: (i) a node can rotate its key pair without the need of refreshing the VC, by updating autonomously the information in the ledger, (ii) a node can immediately revoke its DID if the key pair gets compromised invalidating right away the VC, (iii) the identity life cycle of a device does not require human intervention, and (iv) the scale at which identity can be provided to devices increases dramatically.
@@
Given these benefits, our goal is to contribute to the adoption of SSI in IoT systems and its use for authentication purposes. However, most of the discussions and proposals for the adoption of the SSI model consider an implementation at the application layer of the TCP/IP stack [9, 10]. The problem with application layer authentication is, first, the lack of established protocols that can be used in IoT systems and, second, the need to use two different technologies simultaneously, X.509 certificates on the server side to establish a secure channel and VCs on the client side to complete the mutual authentication on top of the established secure channel. This creates an overall friction in the adoption of the SSI model.
@@
We believe in implementing (mutual) authentication at the transport layer of the TCP/IP stack directly within the Transport Layer Security (TLS) protocol [11] with VC. Using VC in TLS maximizes the benefits of adopting SSI while avoiding the use of PKI and X.509 certificates and their associated costs. In addition, performing authentication at a lower layer in the stack reduces the attack surface and avoids potential vulnerabilities that the application layer may add.
@@
Enabling the use of VCs requires extending the TLS 1.3 handshake protocol to work with VCs in addition to the existing certificate types (e.g., X.509, raw public key) to ensure interoperability. We have already proposed an extension of the TLS 1.3 handshake protocol in [12] enabling the use of VCs and DIDs while preserving all security features of the original handshake. The design has been validated with a real implementation in OpenSSL and statistically relevant experimental results in real-world settings. The design in [12] enables the use of VC and DID for authentication by adding an extension and two new messages, fully compliant with RFC-8446 [11]. Despite full compliance and interoperability our design required a modification to the state machine of the protocol implementation. There is a high cost associated to the modification and associated security review of the TLS handshake state machine. This paper presents a new efficient design of the TLS 1.3 handshake protocol that minimizes the changes with respect to the original one hence make it more appealing for a possible standardization [13] and upstream contribution to the OpenSSL library.
@@
While working on the design and implementation of the new efficient handshake protocol, an interesting work appeared in literature [14]. The authors acknowledge our previous work [12, 15] and they have the same goal of performing authentication in the TLS with VCs and DIDs. Their design divides the overall authentication process in two phases: the pseudo-anonymous authentication with DIDs in the TLS handshake protocol and the identification with Verifiable Presentations (VPs) in the TLS record protocol. However, to leave the handshake unchanged they wrap DIDs inside self-signed X.509 certificates, but this implies the overhead of the certificate generation and management, and the exchange of unnecessary data. In addition, their overall (mutual) authentication process requires a higher number of round trip times (RTTs) than our solutions, because of the identification phase. Furthermore, the pseudo-anonymous authentication with DIDs establish the secure channel even though the identity of the endpoints is still unverified. In a mutual authentication scenario, any adversary can create its own DID and establish the secure channel with the only purpose of draining resources on server side. In our solutions an endpoint never finalizes the handshakes with an unidentified endpoint.
@@
This paper present our new design of an efficient TLS handshake protocol. It does not add any new message or delete any existing one from the original handshake, but only embeds a new extension and adds the VC certificate type to the list of accepted values in the client_certificate_type and server_certificate_type defined in [16]. These two extensions provide by default hybrid handshakes and a fallback mechanism to X.509 certificates in case other certificate types are not supported. This novel approach leaves untouched the core of the handshake protocol, and it is meant to ease adoption. We implemented most of the additional functionalities in an external provider for OpenSSL.
@@ II Self-Sovereign Identity
Layer 1 is implemented by any Distributed Ledger Technology (DLT) acting as a Root of Trust (RoT) for public identity data. DLTs are distributed and immutable means of storage by design [17]. The Decentralized IDentifier (DID) [18] is the new type of globally unique identifier designed to verify a node. The DID is a Uniform Resource Identifier (URI) in the form
@@
While Layer 2 uses DIDs to start authentication, Layer 3 completes it and also deals with authorization to services/resources using Verifiable Credentials (VCs) [20]. A VC is a digital credential that contains additional characteristics of a node’s identity beyond its identity key pair, the DID, and the DID Document.
@@
  • Holder is the node that owns one or more VCs and generates a Verifiable Presentation (VP) to request access to services/resources from a Verifier. A Holder builds the VP as an envelope of the VC and signs it with its identity private key $sk$;
@@ III-A Original handshake
The TLS 1.3 handshake protocol establishes the secure channel by exchanging the messages shown in Fig. 1(a). Upon the handshake, client and server negotiate cryptographic parameters through the exchange of ClientHello and ServerHello messages. Those parameters are the symmetric cipher and hash algorithm used to ensure confidentiality and integrity, respectively. The client and the server generate a secret by using ephemeral Diffie–Hellman (EDH) key exchange and use the HMAC-based Extract-and-Expand Key Derivation Function (HKDF) algorithm with the negotiated hash algorithm to derive the session keys from the secret. Then, the server authenticates with the client by sending the Certificate message containing the certificate chain (Root CA certificate excluded), and the CertificateVerify message which is the signature over all the exchanged messages computed with its private key. Finally, the server sends the Finished message to ensure the integrity of the handshake. It calculates a Hash-Based Message Authentication Code (HMAC) of all the messages exchanged, employing the negotiated hash algorithm, and a Message Authentication Code (MAC) key derived from the session key. The client verifies the validity of the certificate chain, the signature in the CertificateVerify message to check the identity of the server and the HMAC contained in the Finished message. The server can request client authentication by sending the CertificateRequest message. The client authenticates with the server in the same way.
@@ III-B Handshake with VC
We achieve these objectives by leveraging the client_certificate_type and server_certificate_type extensions introduced by RFC-7250 [16] in the TLS 1.3 handshake protocol. These two extensions are used to negotiate certificate types different than X.509 certificates. We added the VC certificate type to the already existing X.509 and RawPublicKey types to enable the use of VCs. The exchange of the messages in the TLS handshake with the new certificate type is represented in Fig.1(b). The client sends the two extensions in the ClientHello indicating the support for the VC certificate type for both client and server authentication. The server sends the two extensions in the EncryptedExtensions message, selecting the VC type from the list presented by the client for both the client and the server.
@@
The did_methods is sent in the ClientHello and EncryptedExtensions messages. A client that sends the server_certificate_type extension containing the VC certificate type must also send the did_methods extension. Similarly, a server that selects VC as the certificate type for the client_certificate_type extension must also send the did_methods extension. A server must send a list of DID Methods both client and server share if it receives the did_methods extension from the client.
@@
The use of VCs reduces the total number of signatures in the handshake. VC Issuers are not organized hierarchically. The server must verify only the signature of the Issuer in the Certificate message. When using X.509 certificates, it must validate the certificate chain up to the Root CA. This is considered an advantage in the case of resource constrained devices, as it can improve performance while maintaining compliance with the VP concept. The server computes the Finished message as in the original TLS 1.3 handshake protocol.
Upon receiving CertificateVerify message, the client checks that the VC follows the scheme specified in the @context field, checks the validity of the VC metadata, resolves the DID of the Issuer and verifies its signature on the VC, and then extracts the server DID from the credentialSubject field of the VC and resolves the server DID to retrieve the server public key from the ledger. Finally, the client verifies the signature in the CertificateVerify message. Note that in the context of IoT systems, the number of available Issuer(s) is expected to be limited (one very often), therefore both client and server could securely maintain the public key of Issuer(s) they trust to avoid the DID resolution. Finally, the client verifies the Finished message. Then, it authenticates with the server in the same way.
@@ IV-A Security
The TLS handshake protocol with VC certificate type maintains the same security features of the original handshake, as the design remains the same. The only addition is the did_methods extension. When the extension is sent in the ClientHello message, it benefits from the integrity and authentication properties. When it is sent in the CertificateRequest message, it benefits also from confidentiality.
@@
Most of the security discussion on our design focuses on the interaction with the DLT node for peer authentication. In the case of a DID resolution performed in clear, an active man-in-the-middle (MiTM) could impersonate the DLT node, forge a DID document containing the authenticating endpoint’s DID, associate it with a key pair that he owns, and then return it to the DID resolver. Thus, the attacker is able to compute a valid CertificateVerify message by possessing the long term private key. In practice, the MiTM attacker breaks in transit the immutability feature of the DLT (i.e., the RoT for identity public keys). To mitigate this risk, an endpoint can establish a secure channel towards the DLT node, either through the Internet Protocol Security (IPSec) [21, 22, 23] or TLS. They both provide authentication of the node and integrity and confidentiality of the data exchanged, but they come with an additional overhead. However, there are some differences among the two approaches.
@@
An IPSec channel works at network layer, and it is established before the TLS handshake takes place and can be reused until it expires. The moment the endpoint performs the DID Resolve, the messages are immediately ciphered and exchanged. An IPSec channel can only be established in case of full control of a DLT node, because additional software must be installed on it. In the case of IoT systems we consider the DLT nodes to be internal to the system and so under full control of the administrator. IPSec employs the Internet Key Exchange (IKE) [22] protocol to authenticate the parties. In IKE authentication must be mutual. One option could be the usage of X.509 certificates, but it is not convenient because all the devices in the system should have an X.509 certificate and so they can directly run the original TLS 1.3 handshake protocol, making VCs worthless. Raw Public Keys (RPKs) [16] represents a better option for authentication in this case. Every IoT node must save locally the public key of each DLT node and each DLT node must save locally the key of each IoT node. The number of DLT nodes within an IoT system is expected to be very low (i.e., one or a couple of nodes) with respect to the total number of IoT and edge nodes in the system, so the number of public keys stored by the latter is minimal. Only the DLT nodes should have the full list of public keys of the devices in the IoT system. One could argue that devices in the IoT system could communicate securely by using directly RPKs for authentication in the TLS handshake, but this means that each IoT and edge node should store locally the public keys of all the other nodes in the system and this is again critical in terms of cost associated to identity management.